Operations Report – April 28th

Crew 193 Operations Report 28Apr2018

SOL: 6

Name of person filing report: David Attig

Non-nominal systems: Generator – sim suit chargers

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 11:00 pm to 8:30 am.

Solar— SOC 70 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 500 gallons

Trailer to Static Pump used – yes

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134250.5

Toilet tank emptied: no
Date last emptied: April 26

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Outreach videos were made today, water transferred from storage tank to static tank. Red wagon assembled for greenhab.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: One piece suits nominal, still work to do on two piece suits. Repair of suit chargers still in progress.

Summary of Hab operations: Transferred water from storage tank to static tank.

Summary of GreenHab operations: Watered twice, plants returning to health. Assembled red wagon placed inside.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: Not used today.

Summary of health and safety issues: One crewmember scraped their hand when going down the stairs (actually fell down the last one).

Questions, concerns and requests to Mission Support: We would like an ice cube tray for the freezer.

Commander's Report – April 28th

Mission Sol 6 (Earth Day April 28th, 2018)

Sols fly by on Mars! Before we knew it, we are at the half-way point in our mission. We took the time today to ponder over what we have accomplished so far and plan the rest of our mission. It being a rest day, we decided to put it to good use and spent the morning creating outreach videos.

Followed it up with a lunch of quinoa, chicken curry and daal.

Our CGA captured a beautiful view of the prominence on the southeast side of the Sun! Cannot wait to take a look at the processed image. Our CE and HSO spent early afternoon putting together the new red wagon as requested by Mission Control.

We plan on taking it easy the rest of the evening. We’ve been learning the American Sign Language from our CGA (who’s hearing impaired), will dedicate some time this evening for the same, a round of "Never Ever Have I" game is on the cards after and might wrap up our evening with a movie, Hidden Figures Or Interstellar. Both interesting. Let’s see what the crew is in mood for.

Thank you for being with us this evening Mission Support!

Signing off for this Sol!

Warm Regards,
Commander, (Gold 00), Crew #193
Mars Desert Research Station

(P.S. Abbreviations Legend: CO – Commander / XO – Executive Officer / CE – Crew Engineer / HSO – Health Science Officer / CGA – Crew Geologist and Astronomer)

Science Report – April 28th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs took place today.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar prominence during his imaging session this morning.

3. Shortwave Texting on EVA’s: No EVAs took place today.

4. GPS Route Measurements: No EVAs took place today.

5. Waterless dish cleaning: Samples were taken from a dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria are growing on the samples from the first test.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No EVAs took place today.

8. Crew Comfort, Health and Safety on EVA: No EVAs took place today.

Sol 6 Summary Report – April 28th

Sol 6

Summary Title: Stopping to Smell the Roses

Author: Eric Shear, CGA

Mission Status: Perfectly nominal… for now.

Sol Activity Summary: This morning we did a series of short outreach videos per the astronomer’s suggestion. For lunch, we feasted on (yet) another delicious Martian lunch of quinoa and curry chicken courtesy of our commander’s mean cooking skills. The afternoon was spent on ongoing science experiments, finishing the MDRS sign, building the new red wagon for the Greenhouse, and usual Hab operations. Our crew astronomer managed to catch a small prominence on the sun’s limb during his observations this morning. The remainder of the day was spent planning out our remaining EVAs. We now have a week’s worth of EVA destinations for both away teams.

Look Ahead Plan: Tomorrow, we will execute the first two EVAs of the new week. In the morning, the American team will head to Zubrin’s Rock and explore White Rock Canyon. Team Canada will follow up at Toothy Ridge in the afternoon.

Anomalies in work:


Weather: Cloudy skies in AM and PM. Extremely hot.

Crew Physical Status: Nominal

EVA: None

Reports filed:

Astronomy Report

GreenHab Report

Reports to be filed:

EVA Report

EVA Request

Science Report

Journalist Report

Photo & Photo of the Day Report

Operations Report

Support Requested:


Astronomy Report – April 28th

Astronomy Report

Name: Eric Shear

Crew: 193
Date: April 28, 2018

Sky Conditions: Sunny with clouds

Wind Conditions: No wind earlier in morning, windy around noon

Observation 1 Start Time: 0945

Observation 1 End Time: 1015

Observation 2 Start Time: 1145

Observation 2 End Time: 1200


Observation 1: The solar flare and filament from yesterday were no longer visible with the camera. No sunspots could be seen with the eye, and the camera could not discriminate any from the granules.

Observation 2: I got a heads-up from Peter Detterline about a prominence on the Sun’s southeast side. It was not visible in the eyepiece at first, but after playing with the tuners it came into view. I attached the Skyris camera, brought the feature into focus, and took a capture.

Photos were taken with the Skyris camera and the best images processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The results are attached.

Objects Viewed: Sun

Problems Encountered: Winds during second observation were such that I could not open the lower shutter without it shaking too much, so I kept it closed. The sun was high enough in the sky by then that I could observe it with just the upper shutter retracted.

GreenHab Report – April 28th

GreenHab Report

Name: Shawna Pandya

Date: 28Apr2018

Environmental control: Ambient with door open

Shade cloth: on

Fan: Used today

Average temperatures: 39.0℃

Low: 15.5℃

High: 52.2℃

Hours of supplemental light: None

Changes to crops: Less yellow with additional watering

Daily water usage for crops: 10 Gallons

Time(s) of watering for crops: Twice: Once at 0900 and again at 1700

Morning research observations: None

Changes to research plants: No.

Aquaponics: None

Narrative: Picked a few tomatoes. Assembled red wagon – the latest ‘Mars Rover’!

Support/supplies needed: None

Copyright © The Mars Society. All rights reserved. | Main Site